OpenStudy (anonymous):

The half life of plutonium is 14.4 years. If 100g of plutonium are present initially, determine the quantity present after 10 years. When will the quantity drop under 5 mg?

OpenStudy (anonymous):

First find the relative decay factor \(k\): \[\frac{1}{2}=e^{14.4k}~~\implies~~k=\frac{1}{14.4}\ln\frac{1}{2}\] Use this \(k\) to solve for the time \(t\): \[5=100e^{kt}~~\implies~~t=\frac{1}{k}\ln\frac{5}{100}\approx\cdots\]